हिंदी

Find the probability that the visitor obtains answer yes from at least 2 pupils: a. when the number of pupils questioned remains at 4. b. when the number of pupils questioned is increased to 8.

Advertisements
Advertisements

प्रश्न

In a large school, 80% of the pupil like Mathematics. A visitor to the school asks each of 4 pupils, chosen at random, whether they like Mathematics.

Find the probability that the visitor obtains answer yes from at least 2 pupils:

  1. when the number of pupils questioned remains at 4.
  2. when the number of pupils questioned is increased to 8.
योग
Advertisements

उत्तर

Let X = number of pupils like Mathematics.

p = probability that pupils like Mathematics

∴ p = 80% = `80/100 = 4/5`

and q = 1 – p = `1 - 4/5 = 1/5`

Given: n = 4

∴ X ~ B `(4, 4/5)`

The p.m.f. of X is given by P(X = x) = `"^nC_x  p^x  q^(n - x)`

i.e. p(x) = `"^4C_x  (4/5)^x  (1/5)^(4 - x)` x = 0, 1, 2, 3, 4

a. P(visitor obtains the answer yes from at least 2 pupils when the number of pupils questioned remains at 4) = P(X ≥ 2)

= P(X = 2) + P(X = 3) + P(X = 4)

`= ""^4C_2  (4/5)^2  (1/5)^(4 - 2) + ""^4C_3  (4/5)^3  (1/5)^(4 - 3) + "^4C_4  (4/5)^4  (1/5)^(4 - 4)`

`= (4 xx 3)/(1 xx 2) xx 16/5^2 xx 1/5^2 + 4 xx 64/5^3 xx 1/5 + 1 xx 256/5^4`

`= 96/5^4 + 256/5^4 + 256/5^4`

`= (96 + 256 + 256)1/5^4`

`= 608/5^4 = 608/625`

b. P(the visitor obtains the answer yes from at least 2 pupils when number of pupils questioned is increased to 8)

= P(X ≥ 2)

= 1 – P(X < 2)

= 1 – [P(X = 0) + P(X = 1)]

= `1 - [""^8C_0  (4/5)^0  (1/5)^(8 - 0) + ""^8C_1  (4/5)^1  (1/5)^(8 - 1)]`

= `1 - [1 (1) (1/5)^8 + (8)(4/5)(1/5)^7]`

= `1 - [1/5^8 + 32/5^8]`

= `1 - 33/5^8`.

shaalaa.com
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 8: Binomial Distribution - Miscellaneous exercise 2 [पृष्ठ २५५]

APPEARS IN

बालभारती Mathematics and Statistics 2 (Arts and Science) [English] Standard 12 Maharashtra State Board
अध्याय 8 Binomial Distribution
Miscellaneous exercise 2 | Q 14.2 | पृष्ठ २५५

वीडियो ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्न

The probability that a certain kind of component will survive a check test is 0.6. Find the probability that exactly two of the next four components tested will survive.


A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of 5 successes. 


A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of at least 5 successes.


Five cards are drawn successively with replacement from a well-shuffled deck of 52 cards; find the probability that all the five cards are spades.


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that exactly one floppy disc work.


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that exactly two floppy disc work.


Choose the correct option from the given alternatives:

For a binomial distribution, n = 5. If P(X = 4) = P(X = 3), then p = ______


For a binomial distribution, n = 4. If 2P(X = 3) = 3P(X = 2), then p = ______.


If X ~ B(4, p) and P(X = 0) = `16/81`, then P(X = 4) = ______.


Let X ~ B(10, 0.2). Find P(X = 1).


Let X ~ B(10, 0.2). Find P(X ≥ 1).


Let X ~ B(10, 0.2). Find P(X ≤ 8).


The probability that a mountain-bike travelling along a certain track will have a tyre burst is 0.05. Find the probability that among 17 riders: exactly one has a burst tyre


The probability that a mountain-bike travelling along a certain track will have a tyre burst is 0.05. Find the probability that among 17 riders: at most three have a burst tyre


An examination consists of 10 multiple choice questions, in each of which a candidate has to deduce which one of five suggested answers is correct. A completely unprepared student guesses each answer completely randomly. What is the probability that this student gets 8 or more questions correct? Draw the appropriate morals.


The probability that a machine will produce all bolts in a production run within specification is 0.998. A sample of 8 machines is taken at random. Calculate the probability that all 8 machines.


The probability that a machine will produce all bolts in a production run within specification is 0.998. A sample of 8 machines is taken at random. Calculate the probability that 7 or 8 machines.


The probability that a machine develops a fault within the first 3 years of use is 0.003. If 40 machines are selected at random, calculate the probability that 38 or more will not develop any faults within the first 3 years of use.


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 1.


In a large school, 80% of the pupil like Mathematics. A visitor to the school asks each of 4 pupils, chosen at random, whether they like Mathematics.
Calculate the probabilities of obtaining an answer yes from 0, 1, 2, 3, 4 of the pupils.


It is observed that it rains on 12 days out of 30 days. Find the probability that it rains exactly 3 days of week.


It is observed that it rains on 12 days out of 30 days. Find the probability that it it will rain at least 2 days of a given week.


In binomial distribution with five Bernoulli’s trials, the probability of one and two success are 0.4096 and 0.2048 respectively. Find the probability of success.


If E(x) > Var(x) then X follows _______.


In Binomial distribution if n is very large and probability success of p is very small such that np = m (constant) then _______ distribution is applied.


In a Binomial distribution with n = 4, if 2P(X = 3) = 3P(X = 2), then value of p is ______.


If X ~ B(n, p) with n = 10, p = 0.4, then find E(X2).


Choose the correct alternative:

A sequence of dichotomous experiments is called a sequence of Bernoulli trials if it satisfies ______


If X follows a binomial distribution with parameters n = 10 and p. If 4P(X = 6) = P(X = 4), then p = ______ 


In a binomial distribution `B(n, p = 1/4)`, if the probability of at least one success is greater than or equal to `9/10`, then n is greater than ______.


In a binomial distribution `B(n, p = 1/4)`, if the probability of at least one success is greater than or equal to `9/10`, then n is greater than ______.


If X∼B (n, p) with n = 10, p = 0.4 then E(X2) = ______.


In a binomial distribution, n = 4 and 2P(X = 3) = 3P(X = 2), then q = ______.


If X is a binomial variable with range {0, 1, 2, 3, 4} and P(X = 3) = 3P(X = 4) then the parameter ‘p’ of the binomial distribution is


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×